# NeuroAMP: A Novel End-to-end General Purpose Deep Neural Amplifier for Personalized Hearing Aids

### v1

@@ 3.1.3 Convolutional Recurrent Neural Network (CRNN)
CRNN combines the strengths of CNN and RNN, making it ideal for signal processing tasks [32, 33] that involve both spatial and temporal feature extraction. Its convolutional layers capture local spectral features, while the recurrent layers model temporal dynamics. This hybrid approach enhances the network’s ability to manage audio signal complexity, thereby improving performance in tasks like speech enhancement [34, 35]. The process of the CRNN in the NeuroAMP model is described as follows:
@@ 3.1.4 Transformer
The Transformer is included for its state-of-the-art performance in sequence modeling tasks [36, 37, 38, 39]. Unlike traditional RNNs, the Transformer uses a self-attention mechanism [36] to process entire sequences in parallel, allowing for efficient handling of long-range dependencies [40]. This capability is particularly beneficial in audio signal processing [41], where capturing the global context and dependencies can enhance performance. The process of the Transformer in the NeuroAMP model is described as follows:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the query, key, and value matrices, $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are the weight matrices for the query, key, and value, $d_{k}$ is the dimension of the key, and FFN is the feedforward network.
@@ 3.1.5 Loss function
To optimize NeuroAMP, we use Mean Squared Error (MSE) as the loss function. This approach penalizes deviations between predicted ($\hat{\textbf{Y}}$) and ground-truth (Y) spectral representations, guiding the network to learn spectral mappings. The loss function is calculated as:
$L_{NeuroAMP}=\frac{1}{T}\sum_{i=1}^{T}|\textbf{y}_{t}-\hat{\textbf{y}}_{t}|^{2},$ (6)
where $\textbf{y}_{t}$ and $\hat{\textbf{y}}_{t}$ are the ground-truth amplified and estimated-amplified spectral representations of the $t$-th frame of the training audio signal $y$, and $T$ is the number of frames.
@@ 3.2 Denoising NeuroAMP Architecture
To address the challenges of noisy environments, which severely impact the effectiveness of hearing aids, we propose the Denoising NeuroAMP system. This system is built on the NeuroAMP architecture, allowing simultaneous denoising and amplification by seamlessly integrating speech enhancement within the neural network. The architecture of the Denoising NeuroAMP model is illustrated in Fig. 3.
@@
As shown in Fig. 3, Denoising NeuroAMP takes a noisy audio signal (X) as input, represented as a magnitude spectrogram. Simultaneously, the listener’s audiogram (Z) is processed through a dense layer to construct a vector representation corresponding to the hearing profile. These two features – the noisy magnitude spectrogram and the audiogram vector – are concatenated and then fed into the core Denoising NeuroAMP neural network. This network, while conceptually similar to the original NeuroAMP model, is trained to perform both speech enhancement and personalized amplification in a single unified process. It learns to adapt its processing based on the characteristics of the noisy input and the specific hearing loss profile, using the audiogram to guide optimization toward the amplified clean audio output.
The training objective of Denoising NeuroAMP, as shown in Eq. 8, is to minimize the MSE between the ground-truth magnitude spectrogram (obtained by applying the NAL-R prescription formula and WDRC to the clean input audio) and the magnitude spectrogram of the predicted enhanced-amplified output from the noisy input. This ensures that the system learns to produce amplified outputs that are both enhanced and aligned with established hearing aid fitting practices.
$\begin{array}[]{c}\mathbf{X}=\text{STFT}(x)\\ \mathbf{Z}=\text{Dense}(z)\\ \mathbf{Concat}=[\mathbf{X}\mid\mathbf{Z}]\\ \mathbf{\tilde{X}}=\text{Den-NeuroAMP}(\mathbf{Concat}).\\ \end{array}$ (7)
$L_{Den-NeuroAMP}=\frac{1}{T}\sum_{i=1}^{T}|\textit{NALR-WDRC}(\textbf{y}_{t})-\tilde{\textbf{x}}_{t}|^{2},$ (8)
The output of the Denoising NeuroAMP network is a predicted magnitude spectrogram of the amplified-enhanced signal. The phase information of the input noisy signal is combined with the predicted magnitude spectrogram to reconstruct the time-domain, enhanced-amplified audio signal through the Inverse Short-Time Fourier Transform (ISTFT).
@@ 4.1 Dataset
In this work, the input to NeuroAMP includes two types of speech signals (English and Taiwanese Mandarin speech data) and music data. We aim to evaluated the proposed models under diverse acoustic properties of cross-linguistic speech and music signals. The speech data contains three conditions: clean, noisy, and enhanced speech. To generate noisy speech, we added 100 types of environmental noise from [42] to clean speech at different signal-to-noise ratio (SNR) levels. Enhanced speech was then obtained by applying two deep learning-based speech enhancement models to the noisy data. The first model is based on LSTM [43], which operates on STFT features to enhance noisy speech. The second model is based on a Fully Convolutional Network (FCN) [27], which operates on the raw waveform to enhance noisy speech. For music data, we mainly consider two categories: clean music and noisy music (i.e., music with added noise). Clean music represents the original recording, while noisy music incorporates various environmental disturbances to simulate real-world listening scenarios. To train the NeuroAMP model, we used TIMIT, TMHINT, and the MUSIC dataset provided by the Cadenza Challenge [44]. To test and evaluate the model’s performance and generalization, we also included the unseen test set (the acoustic conditions and audio content are entirely different from the training conditions) from the VoiceBank and MUSDB18-HQ datasets.
@@ 4.1.1 Training Dataset Description
  • TIMIT Dataset [45]: The TIMIT dataset is a well-known speech corpus used for tasks like speech recognition [46, 47] and speech enhancement [47, 48]. It includes recordings from 630 speakers representing eight major dialects of American English, with each speaker reading 10 phonetically rich sentences. The TIMIT training set consists of 4,620 audio files, while the test set comprises 1,690 audio files. In addition, to augment the dataset, noise signal was injected at different SNR levels(-5 dB, 0 dB, and 5 dB) into the clean audio to create noisy audio. In addition, enhanced versions of these noisy audio were generated using two SE models, which are LSTM and FCN. Following augmentation, a total of 13,700 training samples were selected from the TIMIT dataset, consisting of 3425 clean audio, 3425 noisy audio, and 6850 enhanced audio.
  • TMHINT Dataset: During model training, we incorporated an additional dataset from the Taiwan Mandarin Hearing in Noise Test (TMHINT) to improve the reliability of our models by including diverse acoustic conditions. The TMHINT dataset has been widely used in various speech-related tasks, such as speech assessment [49] and enhancement [47]. Specifically, the TMHINT dataset consists of recordings from multiple speakers, with 16 phonemically balanced sentence sets, each containing 20 sentences. Each sentence comprises ten characters representing diverse and balanced phonetic elements of Mandarin, including four primary lexical tones and a neutral tone. The total number of utterances in this dataset was 1440, which we divided into training and testing sets using an 80/20 ratio, ensuring that speakers in the test set were not used in training. In addition, we expanded the TMHINT dataset by adding environmental noises at different SNR levels (-5 dB, 0 dB, and 5 dB). We also generated enhanced speech from the corresponding noisy speech using LSTM and FCN SE models. We then generated reference data using NAL-R+WDRC, and split this equally between the training and testing sets.
  • MUSIC Dataset: The music dataset from the Cadenza Challenge [44, 50] was incorporated to train the AMP-Net model, utilizing its diverse collection of musical samples. Additionally, to simulate real-world listening conditions, various types of environmental noise were injected into the music data at multiple SNR levels (-5 dB, 0 dB, and 5 dB). This augmentation process was designed to mimic common scenarios where music is played alongside background noise. In total, we selected 11,856 training samples, consisting of 5,928 clean audio samples and 5,928 noisy audio samples.
@@ 4.1.2 Testing Dataset Description
    In addition, we added separate speech and music test sets that were not added to train the model (VoiceBank [51], and MUSDB18-HQ [52]) datasets to evaluate the models’ ability to generalize to the unseen environments. The detailed information for the two unseen datasets is as follows:
  • VoiceBank Dataset [51]: This dataset has been widely employed in various speech-related tasks, such as speech enhancement [53, 54] and recognition [55]. The dataset contains recordings of isolated English digits spoken by multiple speakers, capturing diverse accents and conditions. The test set includes both clean and noisy samples, each containing 824 utterances from 2 speakers. The noisy samples cover four SNR levels: 17.5 dB, 12.5 dB, 7.5 dB, and 2.5 dB, representing different degrees of noise added to the speech signals.
@@ 4.1.3 Denoising NeuroAMP Dataset
For training and evaluating the Denoising NeuroAMP framework, we utilized the VoiceBank+DEMAND dataset [51]. This dataset is widely recognized and employed in the field of speech enhancement, making it particularly suitable for our task of combining denoising with personalized amplification. The VoiceBank+DEMAND dataset comprises the VoiceBank corpus, which contains clean speech recordings from multiple speakers, and the DEMAND corpus, which provides a diverse set of environmental noises. This combination allows the creation of noisy speech samples by mixing clean speech with various types of noise at different SNR levels. The official split of the dataset was used, which includes a training set of 11,572 utterances from 28 speakers and a test set of 824 utterances from 2 speakers. The availability of both clean and noisy speech makes it ideal for training models to perform both speech enhancement and amplification.
@@ 4.1.4 Hearing Loss Patterns
In our study, we use audiograms from the Clarity Challenge, which provides comprehensive samples for both training and testing. These audiograms are essential for evaluating hearing abilities across a standard set of frequencies. We specifically examine audiograms that cover six fundamental frequencies: 250, 500, 1000, 2000, 4000, and 6000 Hz. These frequencies are important for assessing speech intelligibility and understanding the auditory challenges faced by individuals with hearing loss. We follow the approach described by Alshuaib et al. [58] to generate hearing loss patterns, which offers a systematic method for categorizing audiometric data into distinct patterns.
@@ 4.2.1 NeuroAMP Setup
In the NeuroAMP experimental setup, the input signal is preprocessed with a 512-point STFT, a Hamming window of 32 ms, and a hop of 16 ms. The prescription input, which is a 1x6 matrix representing the listener’s audiogram with intensity values for specific frequencies, is processed through dense layers in our model to create a refined representation. This processed audiogram information is then concatenated with the spectral features to form the input vector for further modeling.
The various model architectures and their configurations employed in our study are as follows:
  • Convolutional Neural Network (CNN): The CNN configuration in NeuroAMP includes four convolutional layers with filter sizes of 32, 64, 128, and 256. This setup aims to extract detailed spatial features from the audio signals.
  • Long Short-Term Memory (LSTM): The LSTM model configuration includes two layers, each with 256 units, designed to capture temporal dependencies in the audio signal.
  • Convolutional Recurrent Neural Network (CRNN): The CRNN model combines four convolutional layers with 16, 32, 64, and 128 filter sizes, followed by two LSTM layers, each with 256 units. This structure is particularly effective for handling both spectral features and temporal sequences.
@@
Each model configuration includes a 257-unit dense layer for the final output, ensuring robust feature extraction and signal processing. This architecture enables NeuroAMP to effectively integrate the spectral features from the STFT and the detailed information from the prescription input.
To optimize our model, we have utilized the Adam optimizer with the following hyperparameters: a learning rate of 0.0001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and $\epsilon$ of 1e-07. The training process is configured with a batch size of 1 and 100 epochs. Early stopping is enabled to prevent overfitting, ensuring the system achieves optimal performance.
@@ 4.2.2 Denoising NeuroAMP Setup
For the Denoising NeuroAMP experiments, the setup largely mirrors that of the NeuroAMP, with a few key distinctions. The input signal is similarly preprocessed with a 512-point STFT using a 32 ms Hamming window and a 16 ms hop. The listener’s audiogram, represented as a 1x6 matrix, is processed through a dense layer and concatenated with the spectral features of the noisy input signal. In the Denoising NeuroAMP, we specifically employ an LSTM-based neural network architecture. The LSTM model includes two layers, each with 256 units. This configuration is chosen for its effectiveness in capturing temporal dependencies in the audio signal, which is crucial for both speech enhancement and personalized amplification. The training target for Denoising NeuroAMP is generated by applying the NAL-R prescription formula combined with WDRC to the clean version of the input audio signal. We use the Mean Squared Error (MSE) as the loss function, calculated between the predicted magnitude spectrogram and the target magnitude spectrogram.
Similar to NeuroAMP, the Denoising NeuroAMP model is optimized using the Adam optimizer. The hyperparameters are identical: a learning rate of 0.0001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and $\epsilon$ of 1e-07. The model is trained with a batch size of 1 for 100 epochs, and early stopping is used to prevent overfitting.
@@ 4.3 Evaluation Results
This section outlines the evaluation strategy, focusing on assessing the intelligibility and quality of amplified speech processed by the proposed NeuroAMP model and NAL-R+WDRC (baseline). We employ two established metrics in hearing aid auditory research: HASPI and HASQI for speech, along with the Hearing Aid Audio Quality Index (HAAQI) for music. These metrics are important for quantifying the model’s impact on both speech and music quality. For benchmarking purposes, we compare the metrics derived from NeuroAMP with those obtained from the NAL-R+WDRC processed signal, using three statistical measures: Linear Correlation Coefficient (LCC), Spearman Rank Correlation Coefficient (SRCC), and Mean Squared Error (MSE). Higher LCC values, higher SRCC values, and lower MSE values indicate greater similarity between NeuroAMP and NAL-R+WDRC.
Furthermore, we explore the relationship between model parameter scale and performance through two experimental settings: a basic parameter configuration for all neural networks to establish a baseline (denoted as basic (same)) and an optimized configuration to evaluate the full potential of each architecture (denoted as optimized).

### v2

@@ 3.1.3 Convolutional Recurrent Neural Network (CRNN)
CRNN combines the strengths of CNN and RNN, making it an even better model for signal processing tasks [35, 36] that involve both spatial and temporal feature extraction. Its convolutional layers capture local spectral features, while the recurrent layers model temporal dynamics. This hybrid approach enhances the network’s ability to manage audio signal complexity, thereby improving performance in tasks like speech enhancement [37, 38]. The process of the CRNN in the NeuroAMP model is described as follows:
@@ 3.1.4 Transformer
The Transformer has demonstrated state-of-the-art performance in sequence modeling tasks [39, 40, 41, 42]. Unlike RNNs, Transformer uses a self-attention mechanism [39] to process entire sequences in parallel, allowing efficient handling of long-range dependencies [43]. This capability is particularly beneficial in audio signal processing [44], where capturing the global context and dependencies can enhance performance. Transformer process in the NeuroAMP model is described as follows:
@@
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are the query, key, and value matrices, $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$ are the weight matrices for the query, key and value, $d_{k}$ is the dimension of the key, and FFN represents the feedforward network.
@@ 3.1.5 Loss function
To optimize NeuroAMP, we use Mean Squared Error (MSE) as the loss function. This loss function penalizes deviations between predicted ($\hat{\mathbf{Y}}$) and ground-truth ($\bar{\mathbf{Y}}$) spectral representations, guiding the network to learn spectral mappings. The loss function is calculated as:
$L_{\text{NeuroAMP}}=\frac{1}{T}\sum_{t=1}^{T}\left\|\hat{\mathbf{y}}_{t}-\bar{\mathbf{y}}_{t}\right\|_{2}^{2},$ (6)
where $\bar{\mathbf{y}}_{t}$ and $\hat{\mathbf{y}}_{t}$ are the ground-truth amplified and estimated-amplified spectral representations of the $t$-th frame of the input audio signal $y$, and $T$ is the number of frames.
@@ 3.2 Denoising NeuroAMP Architecture
To address the challenges of noisy environments, which severely impact the effectiveness of hearing aids, we propose the Denoising NeuroAMP system. This system is built on the NeuroAMP architecture, which allows simultaneous denoising and amplification by seamlessly integrating speech enhancement within the NeuroAMP model. The architecture of the Denoising NeuroAMP model is illustrated in Fig. 3.
@@
As shown, Denoising NeuroAMP takes a noisy audio signal (X) as input, represented as a magnitude spectrogram. Simultaneously, the listener’s audiogram is processed through a dense layer to generate a vector representation corresponding to the listener’s hearing profile. These two features – the noisy magnitude spectrogram and the audiogram vector – are concatenated and then fed into the core Denoising NeuroAMP neural network. This network, while conceptually similar to the original NeuroAMP model, is trained to perform speech enhancement and personalized amplification in a single unified process. Based on the noisy input characteristics and the listener’s specific hearing loss profile, Denoising NeuroAMP generates amplified audio.
The training objective of Denoising NeuroAMP, as shown in Eq. 8, is to minimize the MSE between the ground-truth magnitude spectrogram (obtained by applying NAL-R and WDRC to the clean input audio) and the predicted enhanced amplified magnitude spectrogram derived from the noisy input. This ensures that the system learns to produce amplified outputs that are both enhanced and aligned with established hearing aid fitting practices.
[…]
$L_{\text{Den-NeuroAMP}}=\frac{1}{T}\sum_{t=1}^{T}\left|\bar{\mathbf{y}}_{t}-\hat{\mathbf{x}}_{t}\right|^{2},$ (8)
The output of the Denoising NeuroAMP network is a predicted magnitude spectrogram of the amplified enhanced signal. The phase information of the input noisy signal is combined with the predicted magnitude spectrogram to reconstruct the time-domain, enhanced-amplified audio signal through the Inverse Short-Time Fourier Transform (ISTFT).
@@ 4.1 Dataset
In this work, we train the NeuroAMP model by using speech signals of two languages (English and Taiwanese Mandarin speech data) and music data. We aim to evaluate the proposed models under diverse acoustic conditions of cross-linguistic speech and music signals. The speech data contains three conditions: clean, noisy, and enhanced speech. To generate noisy speech, we added 100 types of environmental noise from [45] to clean speech at different signal-to-noise ratio (SNR) levels. Enhanced speech was then obtained by applying two deep learning-based speech enhancement models to the noisy data. The first model is based on LSTM [46], which operates on STFT features to enhance noisy speech. The second model is based on a Fully Convolutional Network (FCN) [30], which operates on the raw waveform to enhance noisy speech. For music data, we mainly consider two categories: clean music and noisy music (i.e., music with added noise). Clean music represents the original recording, while noisy music incorporates various environmental disturbances to simulate real-world listening scenarios. To train the NeuroAMP model, we used TIMIT, TMHINT, and the music dataset provided by the Cadenza Challenge [24]. To test and evaluate the model’s performance and generalization, we also included the unseen test set (the acoustic conditions and audio content are entirely different from the training conditions) from the VoiceBank-DEMAND and MUSDB18-HQ datasets.
@@ 4.1.1 Training Dataset Description
  • TIMIT Dataset: The TIMIT dataset is a well-known speech corpus used for tasks like speech recognition [47, 48] and speech enhancement [48, 49]. It includes recordings from 630 speakers representing eight major dialect regions of American English, with each speaker reading 10 phonetically rich sentences. The TIMIT training set consists of 4,620 audio files, while the test set comprises 1,690 audio files. In addition, to augment the dataset, noise signals were injected at different SNR levels (-5 dB, 0 dB, and 5 dB) into the clean audio to create noisy audio. In addition, Enhanced versions of these noisy audio samples were then generated using two SE models (LSTM and FCN). Following augmentation, a total of 13,700 training samples were prepared based on the TIMIT dataset, consisting of 3,425 clean audio, 3,425 noisy audio, and 6,850 enhanced audio.
  • TMHINT Dataset: The TMHINT dataset was used to improve the reliability of our models by including diverse acoustic conditions. The TMHINT dataset has been widely used in various speech-related tasks, such as speech assessment [50] and enhancement [48]. Specifically, TMHINT dataset consists of recordings from multiple speakers, Each speaker provided 16 phonemically balanced sentence sets, each containing 20 sentences. Each sentence comprises ten characters representing diverse and balanced phonetic elements of Mandarin, including four primary lexical tones and a neutral tone. The total number of utterances in this dataset was 1440, which we divided into training and testing sets using an 80/20 ratio, ensuring that speakers in the test set were not used in training. In addition, we expanded the TMHINT dataset by adding environmental noises at different SNR levels (-5 dB, 0 dB, and 5 dB). We also generated enhanced speech from the noisy speech using LSTM and FCN SE models. We then generated reference data using NAL-R+WDRC, and split this equally between the training and testing sets.
  • MUSIC Dataset: The music dataset from the Cadenza Challenge [24, 51] was used to train the NeuroAMP model. To simulate real-world listening conditions, environmental noise was added at SNR levels of 5-5, 0, and $5$ dB. This augmentation mimics scenarios where music is played with background noise. In total, we used 11,856 training samples: 5,928 clean and 5,928 noisy.
@@ 4.1.2 Testing Dataset Description
    In addition, we included separate speech and music test sets that were not part of the training data (VoiceBank-DEMAND, and MUSDB18-HQ) datasets to evaluate the models’ ability to generalize to the unseen environments. The detailed information for the two unseen datasets is as follows:
  • VoiceBank-DEMAND Dataset: This dataset has been widely employed in various speech-related tasks, such as speech enhancement [52, 53] and recognition [54]. The dataset contains recordings of English speech utterances spoken by multiple speakers, capturing diverse accents and conditions. The test set includes both clean and noisy samples, each containing 824 utterances from 2 speakers. The noisy samples cover four SNR levels: 17.5 dB, 12.5 dB, 7.5 dB, and 2.5 dB, representing different degrees of noise added to the speech signals.
@@ 4.1.3 Denoising NeuroAMP Dataset
We train and evaluate Denoising NeuroAMP on the VoiceBank-DEMAND dataset, which combines clean speech from VoiceBank with diverse environmental noises from DEMAND at various SNRs. Using the official split (11,572 training utterances from 28 speakers; 824 test utterances from 2 speakers), this setup provides clean/noisy pairs ideal for jointly learning speech denoising and personalized amplification.
@@ 4.1.4 Hearing Loss Patterns
In our study, we used audiograms from the Clarity Challenge [19], which provides comprehensive samples for both training and testing. These audiograms are essential for evaluating hearing abilities across a standard set of frequencies. We specifically examine audiograms that cover six fundamental frequencies: 250, 500, 1000, 2000, 4000, and 6000 Hz. These frequencies are important for assessing speech intelligibility and understanding the auditory challenges faced by individuals with hearing loss. We follow the approach described by Alshuaib et al. [57] to generate hearing loss patterns, which offers a systematic method to categorize audiometric data into distinct patterns.
@@ 4.2.1 NeuroAMP Setup
In the NeuroAMP experimental setup, the input signal is preprocessed with a 512-point STFT, a Hamming window of 32 ms, and a hop of 16 ms. The audiogram input, represented as a 1x6 matrix of hearing thresholds at specific frequencies, is processed through dense layers to create a refined representation. This processed audiogram information is then concatenated with the spectral features to form the input vector for further modeling.
The various model architectures and their configurations employed in our study are as follows: [^1] https://github.com/Shafique-Khattak/NeuroAMP
  • CNN: The CNN configuration in NeuroAMP includes four convolutional layers with filter sizes of 32, 64, 128, and 256. This setup aims to extract detailed spatial features from the audio signals.
  • LSTM: The LSTM model configuration includes two layers, each with 256 units, designed to capture temporal dependencies in the audio signal.
  • CRNN: The CRNN model combines four convolutional layers with 16, 32, 64, and 128 filter sizes, followed by two LSTM layers, each with 256 units. This structure is particularly effective for handling both spectral features and temporal sequences.
@@
Each model concludes with a 257-unit dense layer, corresponding to the number of positive frequency bins in a 512-point STFT ($512/2+1=257$). This layer directly predicts the per-bin amplified spectrum for every frame. Because the audiogram embedding is concatenated with the STFT features before entering the neural network, the output is inherently conditioned on the listener’s specific hearing profile. The models were trained for $100$ epochs using the Adam optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-7}$. Training was performed with a batch size of $1$, and early stopping on the validation loss was applied to prevent overfitting.
@@ 4.2.2 Denoising NeuroAMP Setup
For the Denoising NeuroAMP experiments, the setup largely mirrors that of NeuroAMP, with a few key distinctions. Similarly, the input signals are preprocessed using a 512-point STFT, a 32 ms Hamming window, and a 16 ms hop size. The listener’s audiogram, represented as a 1x6 vector, is processed through a dense layer and concatenated with the spectral features of the noisy input signal. In the Denoising NeuroAMP, we specifically employ an LSTM-based neural network architecture. The LSTM model includes two layers, each with 256 units. This configuration is chosen for its effectiveness in capturing temporal dependencies in the audio signal, which is crucial for both speech enhancement and personalized amplification. The training target for Denoising NeuroAMP is generated by applying the NAL-R prescription formula combined with WDRC to the clean version of the input audio signal. We use the MSE as the loss function, calculated between the predicted magnitude spectrogram and the target magnitude spectrogram.
The Denoising NeuroAMP model is optimized using the same Adam optimizer configuration as NeuroAMP (learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-7}$). Training is performed with a batch size of $1$ for $100$ epochs, with an early stop on validation loss to prevent overfitting.
@@ 4.3 Evaluation Results
This section outlines the evaluation strategy, focusing on assessing the intelligibility and quality of amplified speech processed by the proposed NeuroAMP model and NAL-R+WDRC (baseline). We employ two established metrics in hearing aid auditory research: HASPI and HASQI for speech, as well as the HAAQI for music. These metrics are important for quantifying the model’s impact on both speech and music quality. For benchmarking purposes, we compare the metrics derived from NeuroAMP with those obtained from the NAL-R+WDRC processed signal, using three statistical measures: Linear Correlation Coefficient (LCC), SRCC, and MSE. Higher LCC values, higher SRCC values, and lower MSE values indicate greater similarity between NeuroAMP and NAL-R+WDRC.
To provide a robust assessment of model performance, all metrics are reported as mean ± standard deviation. Statistical significance between models was determined using paired t-tests, with p-values below 0.05 considered significant. Significance is denoted in the results tables using appropriate symbols. Furthermore, we explore the relationship between model parameter scale and performance through two experimental settings: a basic parameter configuration for all neural networks to establish a baseline (denoted as basic (same)) and an optimized configuration to evaluate the full potential of each architecture (denoted as optimized).